How many tiles whose length and breadth are 12 cm and 5 cm respectively will be needed to fit in a rectangular region whose length and breadth are respectively?
70 cm and 36 cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Area of rectangle = 70 × 36
= 2520 cm^2
Area of one tile = 5 × 12
= 60 cm^2
Number of tiles = (Area of rectangle) / (Area of one tile)
= 2520 / 60
= 42
Hence, 42 tiles are needed

Thanks